# Global Portraits of Inflation in Nonsingular Variables

### v1

@@ II Scalar-tensor cosmology in metric and Palatini formalisms
$\ddot{\Phi}$ $=-3H\dot{\Phi}-\frac{\mathcal{A}\mathcal{B}^{\prime}+\mathcal{A}^{\prime}\mathcal{B}+3\,\delta_{m}\mathcal{A}^{\prime}\mathcal{A}^{\prime\prime}}{2\mathcal{A}\mathcal{B}+3\,\delta_{P}(\mathcal{A}^{\prime})^{2}}\dot{\Phi}^{2}-\frac{2\mathcal{A}\mathcal{V}^{\prime}-4\mathcal{V}\mathcal{A}^{\prime}}{2\mathcal{A}\mathcal{B}+3\,\delta_{m}(\mathcal{A}^{\prime})^{2}}\,.$ (5c)
@@
$\mathcal{B}_{\rm Palatini}$ $=\mathcal{B}_{\rm metric}+\frac{3(\mathcal{A})^{\prime}}{2\mathcal{A}}\,.$ (11)
@@
In the case of minimal coupling ($\mathcal{A}\equiv 1$), $\mathrm{w}_{\mathrm{eff}}$ varies between $-1$ (occurring in the potential dominated regime where the scalar field derivatives can be neglected over the potential), and $+1$ (occurring in the kinetic regime where the potential can be neglected over the derivatives). In the case of nonminimal coupling, the effective barotropic index can take a wider range of values, including the superaccelerating regime where $\mathrm{w}_{\mathrm{eff}}<-1$ and $\dot{H}>0$, as well as superstiff regime where $\mathrm{w}_{\mathrm{eff}}>1$. When the value of $\mathrm{w}_{\mathrm{eff}}$ is constant, we can integrate the left side of (15) to get
@@
where $t_{0}$ and $a_{0}$ are integration constants. The scalar field finite fixed points are always de Sitter (16b) [38, 45, 47, 32, 51, 65, 72]. But for the global portraits it is important to know that the scalar field asymptotic states can correspond to power law expansion (16a) with different values of $\mathrm{w}_{\mathrm{eff}}$ [32, 26, 31], a de Sitter state (16b) [32, 33, 26], or also an “asymptotic de Sitter” state whereby $H$ and $|\Phi|$ diverge, but $\mathrm{w}_{\mathrm{eff}}$ as defined in (15) is still $-1$ [34, 26].
@@
Physical situations correspond to real $H$ while in the context of inflation we are interested in expanding universe where $H>0$. In fact, our assumptions (2), (3) already guarantee that $H$ is real for all values of $\Phi$ and $\dot{\Phi}$, due to (18b). The expression (18a) tells that in the metric case the ‘$+$’ branch corresponds to an expanding universe provided $\mathcal{B}>0$, but in the Palatini case the ‘$+$’ branch may involve a contracting behavior well, for instance when $\dot{\Phi}\mathcal{A}^{\prime}>0$ and $\mathcal{V}$ is small, i.e. during the late phase after inflation where the scalar field oscillates around the minimum of the potential.
@@ III Inflation as a dynamical system
While all values of the original quantities $\Phi$ and $\dot{\Phi}$ corresponded to two real values of $H$, it is not guaranteed that rewriting the constraint (18) in terms of the dynamical variables (19) yields real $H$ for the whole span of the new variables. Obviously, the part of phase space where $\phi$ and $z$ yield complex $H$ is not physical. Another complication stems from the fact that $H$ in (18) is double valued, this is the nature of the Friedmann constraint independent of whether it is expressed in terms of $\Phi$ and $\dot{\Phi}$, or $\phi$ and $z$. Therefore, it requires extra attention to keep track how the “$+$” and “--” sheets of $H(\Phi,\dot{\Phi})$ are mapped to the “$+$” and “--” sheets of $H(\phi,z)$. The relationship can become rather tricky when $z$ is defined to involve $H$ itself. In the present study when drawing the global portraits we plotted the “$+$” sheet of $H(\phi,z)$, calling physical the region where $H$ is real, matching in value to $H(\Phi,\dot{\Phi})$, and which is contiguous to the slow roll inflationary trajectory. Otherwise, especially in the Palatini case, the technically complete analysis would have introduced extra mathematical details that are irrelevant for the physics of inflation, and do not contribute to the main import of the paper. With these premises and as a consistency check we made sure that the boundary of the physical phase space is a trajectory itself that blocks other trajectories to cross over between the physical and nonphysical parts of the phase space.
@@ IV.2 Hubble rescaled variables in e-folds time
$\zeta_{\pm}$ =±$=\frac{\sqrt{\pm 12\sqrt{6}\xi^{\frac{3}{2}}\sqrt{6\xi+1}+108\xi^{2}+6\xi}}{\sqrt{180\xi^{2}+12\xi+1}}\,.$ (38)
@@
$A$ $(0,0)$ nonhyperbolic $-1-8\xi$
$C_{\pm}$ (±$\left(\pm 1,0\right)$ nonhyperbolic $-1$
$D_{\pm}$ (±$\left(\pm\sqrt{1-\zeta_{-}^{2}},\mp\zeta_{-}\right)$ unstable node $1+8\xi-\frac{4\sqrt{6\xi+36\xi^{2}}}{3}$
$E_{\pm}$ (±$\left(\pm\sqrt{1-\zeta_{+}^{2}},\pm\zeta_{+}\right)$ saddle $1+8\xi+\frac{4\sqrt{6\xi+36\xi^{2}}}{3}$
@@ V.1 Direct scalar field variables in e-folds time
Thus there are two types of fixed points on the asymptotic circle, at (±$(\pm 1,0)$ and $(0,\pm 1)$. Their characteristic matrices diverge, but we can still identify the first points with $C_{\pm}$ and the second points with a merger of $D_{\pm}$ and $E_{\pm}$ of Table 1. In the latter case the effective barotropic index depends on the direction of approach. Indeed, we can compute the limits to find
@@ V.4 Hybrid variables with adjustable rescaling in hybrid time
For large $H$ of the early universe the $s$ parameter plays no role and leaves the asymptotic initial states as before. For small $H$ of the late universe the $s$ parameter makes the modified hybrid variable $\hat{z}$ bigger than the previous hybrid variable $\tilde{z}$ when $s<1$ and smaller if $s>1$. Therefore depending on the situation the inspiral into the final point $A$ can be clearly depicted by choosing a suitable value for $s$. Since $s$ is positive, no new fixed points are introduced, and neither will it change the properties of the fixed points from the ones known previously in the hybrid variables. Its only effect is to “stretch out” or “squash” the vertical axis near the point $A$, and enable the optimal visual presentation of the dynamics.
@@ VI.4 Nonminimal Palatini model
The corresponding global phase portrait is given on Fig. 4. In terms of the fixed point structure, unphysical regions, and generic phase flow the diagram is qualitatively similar to the metric Higgs model on Fig. 3 and 4, while the slow roll dynamics is pushed to take place at very high values of the scalar field. The main difference is that in the Palatini case the “$+$” sheet of $H(\phi,\hat{z})$ that we plot here, can also feature contracting behavior of the universe, as explained after Eq. (18) and again before Eq. (25). On the diagram 4 the contracting phase is depiced by the semi-transparent blue shade, and we can notice how trajectories enter the contracting phase in the superstiff ($\mathrm{w}_{\mathrm{eff}}>1$, yellow) region and exit back into expansion in the superaccelerated ($\mathrm{w}_{\mathrm{eff}}<-1$, dark green) region. In fact, the phase portrait tells that in approaching the late universe all trajectories will periodically experience alternating contracting and expanding phases. Thus, unless the effects of reheating drastically change this feature, the current model does probably not offer a realistic scenario to describe our Universe.
Our plot 4 of the nonminimal Palatini model (55) in the hybrid variables (49) and hybrid time (47) can be compared to Fig. 2d of Ref. [31] which depicts the same model with the same parameter values, but just using the Hubble rescaled variables (33) and e-folds time (20). Besides the obvious aspect that the latter is singular at the origin, an attentive reader may also notice how the plot in [31] features two extra “pockets” or regions full of trajectories in the phase space. Although these areas still belong to the real part of the “$+$” branch of the Friedmann constraint $H(\phi,\hat{z})$, in the present work we have excluded those regions from the physically interesting phase space, since they feature contracting behaviour and are discontinuous from the region which inhabits the slow roll trajectory. The solutions in those “pockets” do not enjoy the privilege of being attracted to the slow roll inflationary regime, and thus are not really relevant from the physical point of view.

### v2

@@ II Scalar-tensor cosmology in metric and Palatini formalisms
$\ddot{\Phi}$ $=-3H\dot{\Phi}-\frac{\mathcal{A}\mathcal{B}^{\prime}+\mathcal{A}^{\prime}\mathcal{B}+3\,\delta_{m}\mathcal{A}^{\prime}\mathcal{A}^{\prime\prime}}{2\mathcal{A}\mathcal{B}+3\,\delta_{m}(\mathcal{A}^{\prime})^{2}}\dot{\Phi}^{2}-\frac{2\mathcal{A}\mathcal{V}^{\prime}-4\mathcal{V}\mathcal{A}^{\prime}}{2\mathcal{A}\mathcal{B}+3\,\delta_{m}(\mathcal{A}^{\prime})^{2}}\,.$ (5c)
@@
$\mathcal{B}_{\rm Palatini}$ $=\mathcal{B}_{\rm metric}+\frac{3(\mathcal{A}^{\prime})^{2}}{2\mathcal{A}}\,.$ (11)
@@
In the case of minimal coupling ($\mathcal{A}\equiv 1$), $\mathrm{w}_{\mathrm{eff}}$ varies between 1-1 (occurring in the potential dominated regime where the scalar field derivatives can be neglected over the potential), and +1+1 (occurring in the kinetic regime where the potential can be neglected over the derivatives). In the case of nonminimal coupling, the effective barotropic index can take a wider range of values, including the superaccelerating regime where $\mathrm{w}_{\mathrm{eff}}<-1$ and $\dot{H}>0$, as well as superstiff regime where $\mathrm{w}_{\mathrm{eff}}>1$. When the value of $\mathrm{w}_{\mathrm{eff}}$ is constant, we can integrate the left side of (15) to get
@@
where $t_{0}$ and $a_{0}$ are integration constants. The scalar field finite fixed points are always de Sitter (16b) [38, 45, 47, 32, 51, 65, 72]. But for the global portraits it is important to know that the scalar field asymptotic states can correspond to power law expansion (16a) with different values of $\mathrm{w}_{\mathrm{eff}}$ [32, 26, 31], a de Sitter state (16b) [32, 33, 26], or also an “asymptotic de Sitter” state whereby $H$ and $|\Phi|$ diverge, but $\mathrm{w}_{\mathrm{eff}}$ as defined in (15) is still 1-1 [34, 26].
@@
Physical situations correspond to real $H$ while in the context of inflation we are interested in expanding universe where $H>0$. In fact, our assumptions (2), (3) already guarantee that $H$ is real for all values of $\Phi$ and $\dot{\Phi}$, due to (18b). The expression (18a) tells that in the metric case the ‘++’ branch corresponds to an expanding universe provided $\mathcal{B}>0$, but in the Palatini case the ‘++’ branch may involve a contracting behavior well, for instance when $\dot{\Phi}\mathcal{A}^{\prime}>0$ and $\mathcal{V}$ is small, i.e. during the late phase after inflation where the scalar field oscillates around the minimum of the potential.
@@ III Inflation as a dynamical system
While all values of the original quantities $\Phi$ and $\dot{\Phi}$ corresponded to two real values of $H$, it is not guaranteed that rewriting the constraint (18) in terms of the dynamical variables (19) yields real $H$ for the whole span of the new variables. Obviously, the part of phase space where $\phi$ and $z$ yield complex $H$ is not physical. Another complication stems from the fact that $H$ in (18) is double valued, this is the nature of the Friedmann constraint independent of whether it is expressed in terms of $\Phi$ and $\dot{\Phi}$, or $\phi$ and $z$. Therefore, it requires extra attention to keep track how the “++” and “-” sheets of $H(\Phi,\dot{\Phi})$ are mapped to the “++” and “-” sheets of $H(\phi,z)$. The relationship can become rather tricky when $z$ is defined to involve $H$ itself. In the present study when drawing the global portraits we plotted the “++” sheet of $H(\phi,z)$, calling physical the region where $H$ is real, matching in value to $H(\Phi,\dot{\Phi})$, and which is contiguous to the slow roll inflationary trajectory. Otherwise, especially in the Palatini case, the technically complete analysis would have introduced extra mathematical details that are irrelevant for the physics of inflation, and do not contribute to the main import of the paper. With these premises and as a consistency check we made sure that the boundary of the physical phase space is a trajectory itself that blocks other trajectories to cross over between the physical and nonphysical parts of the phase space.
@@ IV.2 Hubble rescaled variables in e-folds time
$\zeta_{\pm}$ $=\frac{\sqrt{\pm 12\sqrt{6}\xi^{\frac{3}{2}}\sqrt{6\xi+1}+108\xi^{2}+6\xi}}{\sqrt{180\xi^{2}+12\xi+1}}\,.$ (38)
@@
AA $(0,0)$ nonhyperbolic $1-8\xi$
$C_{\pm}$ $\left(\pm 1,0\right)$ nonhyperbolic 1-1
$D_{\pm}$ $\left(\pm\sqrt{1-\zeta_{-}^{2}},\mp\zeta_{-}\right)$ unstable node $1+8\xi-\frac{4\sqrt{6\xi+36\xi^{2}}}{3}$
$E_{\pm}$ $\left(\pm\sqrt{1-\zeta_{+}^{2}},\pm\zeta_{+}\right)$ saddle $1+8\xi+\frac{4\sqrt{6\xi+36\xi^{2}}}{3}$
@@ V.1 Direct scalar field variables in e-folds time
Thus there are two types of fixed points on the asymptotic circle, at $(\pm 1,0)$ and $(0,\pm 1)$. Their characteristic matrices diverge, but we can still identify the first points with $C_{\pm}$ and the second points with a merger of $D_{\pm}$ and $E_{\pm}$ of Table 1. In the latter case the effective barotropic index depends on the direction of approach. Indeed, we can compute the limits to find
@@ V.4 Hybrid variables with adjustable rescaling in hybrid time
For large $H$ of the early universe the $s$ parameter plays no role and leaves the asymptotic initial states as before. For small $H$ of the late universe the $s$ parameter makes the modified hybrid variable $\hat{z}$ bigger than the previous hybrid variable $\tilde{z}$ when $s<1$ and smaller if $s>1$. Therefore depending on the situation the inspiral into the final point $A$ can be clearly depicted by choosing a suitable value for $s$. Since $s$ is positive, no new fixed points are introduced, and neither will it change the properties of the fixed points from the ones known previously in the hybrid variables. Its only effect is to “stretch out” or “squash” the vertical axis near the point $A$, and enable the optimal visual presentation of the dynamics. While the dimensionful parameter $M$ still completely cancels out of the dynamical system, the dimensionless parameter $s$ enters the equations. It does not introduce a new overall energy scale into the original system, though, but scales the dimensionless speed variable $\hat{z}$ in comparison to the dimensionless field variable $\phi$ by increasing/decreasing their ratio as an artificial “slow/fast motion” device in the depiction of the late stages of the evolution.
@@ VI.4 Nonminimal Palatini model
[…]